OpenStudy (krispy):

Solve for p. 3+|1+p|=12 A.p = 8 or p=−10 B.p = 14 or p=−16 C.p = 10 or p=−10 D.p = 8 or p=−8

Mehek (mehek14):

First we need to get rid of 3 Subtract 3 on both sides |1 + p| = 9 Now we need 2 equations 1 + p = 9 1 + p = - 9 Solve for p in both

OpenStudy (triciaal):

here are the steps isolate the term with the subject you need undo the modulus so you have 2 equations solve for the variable/subject always maintain the balance of an equation

OpenStudy (krispy):

Ya'know this is why PA is better because they give the answer instead of telling you how to do it WITHOUT the answer.

Mehek (mehek14):

If you are told how to solve it, you won't have to depend on others to solve it for you.

OpenStudy (krispy):

I'm to dumb to look at that.

Mehek (mehek14):

It's okie ^.^ You're not dumb (:

OpenStudy (krispy):

Then why'd you answer the question

Mehek (mehek14):

Because you asked it?